OpenStudy (anonymous):

Find the equation of the line that contains the point (-4, 18) and is parallel to the line that contains (2, -2) and (6, 3) y=mx+b form, do not use decimals. Leave improper fractions improper. help?

OpenStudy (anonymous):

x1=2 , y1=-2 x2=6 , y2 = 3 the eqn of line passing through (x1,y1) and (x2,y2) can be written as ; (y2-y1)/(x2-x1) = (y-y1)/(x-x1) substituting (x1,y1) and (x2,y2) in above eqn we get ; 5/4 = (y+2)/(x-2) which gives us , y=(5/4)*x - (9/2) any line parallel to above line must have same slope and can be written as y=(5/4)*x + c ( 'c' has to be determined ) this line passes through (-4,18) hence x=-4 and y = 18 must satisfy it 18 = (5/4)*(-4) + c therefore c = 23 hence the required eqn is , y = (5/4)*x + 23

OpenStudy (anonymous):

O wow, thank you!

OpenStudy (anonymous):

ok its telling me ths is wrong, i guess im not distributing the 5/4 properly